### v1

@@
Abstract    Smoluchowski-Kramers approximation in rough path topology with state-dependent damping is explored. The second-order Langevin equation has a form of fast-slow system after suitable change-of-variable, and then its solution is lifted as a rough path in a natural manner. Moment estimates of both the original path and the lift are given, followed by which, averaging technique and convergence theorem in rough path topology are used to pass the limit.
Keywords    Smoluchowski–Kramers approximation, Averaging, Singular pertubation, Rough path theory
@@ 1 Introduction
Consider the state-dependent second-order Langevin equation
$\epsilon^{2}\ddot{X}^{\epsilon}_{t}+M(X^{\epsilon}_{t})\dot{X}^{\epsilon}_{t}=F(X^{\epsilon}_{t})+\dot{W}_{t},\kern 5.0ptX^{\epsilon}_{0}=0,\kern 5.0pt\dot{X}^{\epsilon}_{0}=0,$ (1.1)
defined on a complete stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},P),$ where $0<T<\infty$ is fixed, $M$ is a matrix-valued function and $F$ is a non-linearity, detailed in Section 2. We are interested in the limit of $X^{\epsilon}$ as $\epsilon\to 0,$ which is called Smoluchowski–Kramers (SK) approximation. There is plenty of research for the case $M$ is a constant [2, 3, 33, 16, 30, 7], and in this case, it is not surprising that $X^{\epsilon}\to X$ in some sense with
$M\dot{X}_{t}=F(X_{t})+\dot{W}_{t},\kern 5.0ptX_{0}=0,$
which can be obtained formally by taking $\epsilon=0$ in (1.1). However, when $M$ is not a constant, that is, it is state-dependent, the limit cannot be obtained by simply dropping the small mass term. In fact, under certain assumptions, Hottovy et al.[12] showed that
@@
$M(x)J(x)+J(x)M^{\top}(x)=\mathrm{id}.$ (1.3)
Since then, there are fruitful work on SK approximation with state-dependent damping. For example, Wang et al.[27] explore the limit by averaging on the Fokker-Planck equation associated with the Langevin equation. Cerrai and Xi[4] consider the infinite dimensional case. For more research in SK approximation in state dependent case, we refer to [5, 11, 17, 28].
SK approximation is naturally linked to stochastic averaging, initiated by Khaminskii [13]. Numerous work concerning about stochastic averaging appeared in recent years, see e.g. [14, 29, 32, 31]. There is also research about averaging principle in rough path framework, see for example in [15, 25]. In fact, by suitable change-of-variable, (1.1) is transformed to (2.3), which has a form of fast-slow system [6], making it possible to find its limit by averaging method. A Poisson equation technique[21, 22, 23] is employed to show the convergence.
Rough path theory is initially proposed by Lyons[19]. Although a deterministic theory, it is proved a powerful tool in stochastic analysis. See monographs [18, 20, 9, 8] for systematic display of rough path theory. In this paper, we aim to find the small mass limit in rough path topology. Friz et al [7] identify the limit in rough path topology in the case that $M$ is a constant matrix and $F\equiv 0$. They found that, the limit of the "level-2" rough path, converges to a non-trivial limit. In particular, when the dimension is $1$, the Itô lift of the original path, converges to the Stratonovich lift of the limit path, which is reminiscent of Wong–Zakai approximation [26]. They also gave a conjecture for the case when $M$ is not constant, without detailed proof. In this paper, the system with both state-dependent friction and non-linearity is explored. We identify the limit in the rough path topology, and give rigorous proof. This also generalizes the result in [12], in the sense that in our main theorem, the "level-1" path converges in $p$-th moment of $\alpha$-Hölder norm for all $1\leq p<\infty$.
@@
The rest of this paper is organized as follows. In Section 2, we recall some basic rough path theory, list some notations, and present our main result. In Section 3, we give moment estimate in rough path norm. In Section 4, we display an averaging principle for specific fast-slow system, which is used to pass the limit. After these preparation, we prove the main theorem in Section 5.
2 Preliminary and Main Result
@@
We recall some basic definitions of rough path theory [8, e.g.]. Let $X:[0,T]\to\mathbb{R}^{d},$ $\mathbb{X}:\Delta_{[0,T]}\to\mathbb{R}^{d\times d}$, where
$\Delta_{[0,T]}:=\{(s,t)\in[0,T]^{2}:s<t\}.$
@@
(i) $X||_{\alpha}:=\sup_{0\leq s\leq t}\limits\frac{|X_{t}-X_{s}|}{|t-s|^{\alpha}}<\infty$
(ii) $\mathbb{X}||_{2\alpha}:=\sup_{0\leq s\leq t}\limits\frac{|\mathbb{X}_{s,t}|}{|t-s|^{2\alpha}}<\infty$
(iii) (Chen relation) $\mathbb{X}_{s,u}=\mathbb{X}_{s,t}+\mathbb{X}_{t,u}+X_{s,t}\otimes X_{t,u},\kern 5.0pt\forall 0\leq s\leq t\leq u\leq T.$
The set of all $\alpha$-Hölder rough path over $\mathbb{R}^{d}$ is denoted as $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}).$ Given $\mathbf{X},\mathbf{Y}\in\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}),$ the $\alpha$-Hölder rough path metric is defined as
@@
Let $V^{\epsilon}_{t}=\dot{X}^{\epsilon}_{t}$ and $P^{\epsilon}_{t}=\epsilon^{2}V^{\epsilon}_{t}$ be velocity and momentum, respectively, then (1.1) is equivalent to
@@
Here we also introduce $Y^{\epsilon}_{t}:=\frac{1}{\epsilon}P^{\epsilon}_{t}$ [7, Page 7944], so that
For each $0<\epsilon\leq 1,$ define the lift of $X^{\epsilon}$ by
$\mathbb{X}^{\epsilon}_{s,t}:=\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u},\quad(s,t)\in\Delta_{[0,T]},$
where the stochastic integral is Itô integral, and set $\mathbf{X}^{\epsilon}=(X^{\epsilon},\mathbb{X}^{\epsilon}).$ It is straightforward to check that $\mathbf{X}^{\epsilon}$ satisfies Chen relation. Moreover, by Kolmogorov criterion for rough paths [8, Theorem 3.1] and Proposition 3.3 below, $\mathbf{X}^{\epsilon}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
@@
Now we state the main result.
@@
By Kolmogorov criterion for rough paths again, one verifies that $\mathbf{X}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
@@
$A^{\top}$: Transpose of a matrix or a vector $A$.
$\mathrm{id}$: Identity matrix.
@@
$C^{k}(U;V)$: The set consists of $k$-th continuously differentiable functions from $U$ to $V$, where $k\in\mathbb{N},$ and $U$, $V$ are Banach spaces.
@@ 3 Moment Estimates
In this section, we give several moment estimate for system (2.3). We begin with bounding the supremum norm.
@@
$\sup\limits_{0<\epsilon\leq 1}\sup\limits_{0\leq t\leq T}\mathbb{E}|Y^{\epsilon}_{t}|^{p}<\infty,\kern 5.0pt\sup\limits_{0<\epsilon\leq 1}\mathbb{E}\sup\limits_{0\leq t\leq T}|X^{\epsilon}_{t}|^{p}<\infty.$
@@
We firstly show the boundedness for $Y^{\epsilon}$. For each $p\geq 2,$ set $\phi(y)=|y|^{p},$ then
$D\phi(y)=p|y|^{p-2}y,\quad D^{2}\phi(y)=p(p-2)|y|^{p-4}yy^{\top}+p|y|^{p-2}\mathrm{id}.$
Apply Itô formula to $\phi(Y^{\epsilon}_{t})$ and use ($\mathbf{A1}$) and ($\mathbf{A4}$) to find that
@@
$+\frac{1}{2\epsilon^{2}}\int_{0}^{t}\mathrm{Tr}[p(p-2)|Y^{\epsilon}_{s}|^{p-4}Y^{\epsilon}_{s}(Y^{\epsilon}_{s})^{\top}+p|Y^{\epsilon}_{s}|^{p-2}\mathrm{id}]$
$\lesssim$ $-\frac{1}{\epsilon^{2}}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p}ds+\frac{1}{\epsilon}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-1}ds+H^{\epsilon}_{t}+\frac{1}{\epsilon^{2}}\int_{0}^{t}|Y^{\epsilon}_{s}|^{p-2}ds,$
where $H^{\epsilon}$ is a martingale. Taking expectation and applying Hölder inequality,
$\mathbb{E}|Y^{\epsilon}_{t}|^{p}\lesssim-\frac{1}{\epsilon^{2}}\int_{0}^{t}\mathbb{E}|Y^{\epsilon}_{s}|^{p}ds+\frac{1}{\epsilon^{2}}\int_{0}^{t}\mathbb{E}|Y^{\epsilon}_{s}|^{p-2}ds.$ (3.2)
@@
Next, we show the boundedness for $X^{\epsilon}$. Consider the following random ODE in $\mathbb{R}^{d\times d}$
$\dot{\phi}^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})\phi^{\epsilon}_{t},\kern 5.0pt\phi^{\epsilon}_{0}=\mathrm{id}.$
For each $r\in[0,T],$ set $z^{\epsilon,r}_{s}=\phi^{\epsilon}_{s}(\phi^{\epsilon}_{r})^{-1},\kern 5.0pts\in[r,T].$ It is showed in [27, Page 6] that, under assumptions ($\mathbf{A1}$)–($\mathbf{A4}$),
$X^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}dW_{r}ds+\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})drds.$ (3.3)
and that
$|z^{\epsilon,r}_{s}|\leq e^{-\epsilon^{-2}\lambda(s-r)},\quad 0\leq r\leq s\leq T.$ (3.4)
For notation simplicity, we always assume that $\lambda=1$. Then, from (3.4)
$\Big|\int_{r}^{t}z^{\epsilon,r}_{s}ds\Big|\leq\int_{r}^{t}|z^{\epsilon,r}_{s}|ds\leq\epsilon^{2}.$ (3.5)
Notice that by stochastic Fubini theorem, we have
$X^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}dsdW_{r}+\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr.$ (3.6)
An application of Burkholder–Davis–Gundy inequality and (3.5) yield
$\frac{1}{\epsilon^{2p}}\mathbb{E}\sup\limits_{0\leq t\leq T}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}dsdW_{r}\Big|^{p}$ (3.7)
$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{0}^{T}\Big|\int_{r}^{t}z^{\epsilon,r}_{s}ds\Big|^{2}dr\Big)^{p/2}$
$\lesssim$ $1.$
By ($\mathbf{A4}$) and (3.5), for each $t\in[0,T],$
$\frac{1}{\epsilon^{2p}}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr\Big|^{p}$ (3.8)
$\leq$ $\frac{1}{\epsilon^{2p}}\Big(\int_{0}^{t}\int_{r}^{t}|z^{\epsilon,r}_{s}|dsdr\Big)^{p}$
$\leq$ $T^{p}.$
Taking supremum and expectation,
$\frac{1}{\epsilon^{2p}}\mathbb{E}\sup\limits_{0\leq t\leq T}\Big|\int_{0}^{t}\int_{r}^{t}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})dsdr\Big|^{p}\lesssim 1.$ (3.9)
The result follows from combining (3.6), (3.7) and (3.9). ∎
We need the following elementary inequality, which is a special case of Minkowski integral inequality[10, Page 13]. For completeness, we give its proof.
Lemma 3.2.
Let $G$ be a random process on $[0,T],$ $a,b]\subset[0,T],$ $q\in\mathbb{N},$ then
$\Big[\mathbb{E}\Big(\int_{a}^{b}|G(r)|^{2}dr\Big)^{q}\Big]^{1/q}\leq\int_{a}^{b}\Big(\mathbb{E}|G(r)|^{2q}\Big)^{1/q}dr.$
Proof.
Applying Hölder inequality,
$\mathbb{E}\Big(\int_{a}^{b}|G(r)|^{2}dr\Big)^{q}$
$=$ $\mathbb{E}\int_{[a,b]^{q}}|G(r_{1})|^{2}\cdots|G(r_{q})|^{2}dr_{1}\cdots dr_{q}$
$=$ $\int_{[a,b]^{q}}\mathbb{E}(|G(r_{1})|^{2}\cdots|G(r_{q})|^{2})dr_{1}\cdots dr_{q}$
$\leq$ $\int_{[a,b]^{q}}(\mathbb{E}|G(r_{1})|^{2q})^{1/q}\cdots(\mathbb{E}|G(r_{q})|^{2q})^{1/q}dr_{1}\cdots dr_{q}$
$=$ $\Big(\int_{a}^{b}(\mathbb{E}|G(r)|^{2q})^{1/q}dr\Big)^{q}.$
@@
In view of Jensen inequality, it is sufficient to prove the proposition for all even $p.$ We first prove (i). From (3.3) and stochastic Fubini theorem,

### v2

@@
Abstract    We explore the limit of stochastic differential equations driven by some random processes satisfying singularly perturbed second order stochastic differential equations. The main tool we employ is the universal limit theorem in rough path theory. To this end, we lift the random process as a rough path in a natural manner. After suitable change-of-variable, the random process has a form of slow-fast system. Moment estimates of both the random process and its lift are given, followed by which, averaging technique and convergence theorem in rough path topology are used to identify the limit.
Keywords    Smoluchowski–Kramers approximation, Averaging, Singular perturbation, Rough path theory
@@ 1 Introduction
Let us consider approximation of stochastic differential equations (SDE) of the form
$dZ^{\epsilon}_{t}=b(Z^{\epsilon}_{t})dt+\sigma(Z^{\epsilon}_{t})dX^{\epsilon}_{t},$ (1.1)
where $X^{\epsilon}$ is some random process converging to $X$ in some sense. At first glance, one might expect that $Z^{\epsilon}\to Z$ with
$dZ_{t}=b(Z_{t})dt+\sigma(Z_{t})dX_{t}.$
However, this is not true in general, as seen in the classical Wong–Zakai approximation, and there is an additional term called Itô–Stratonovich correction, appearing in the limit equation[26].
In the pioneering work [7], they assumed that the process $X^{\epsilon}$ is governed via the following Langevin equation
$\epsilon^{2}\ddot{X}^{\epsilon}_{t}+M\dot{X}^{\epsilon}_{t}=\dot{W}_{t},\quad X^{\epsilon}_{0}=0,\kern 5.0pt\dot{X}^{\epsilon}_{0}=0.$ (1.2)
The process $X^{\epsilon}$ is also called physical Brownian motion (pBm). This equation describes the motion of a particle with mass $\epsilon^{2}$ under a random external force $W$, classically assumed to be a standard (mathematical) Brownian motion, and $M$ is the coefficient of friction. They firstly showed that $X^{\epsilon}\to X:=M^{-1}W$, which can be formally obtained by setting $\epsilon=0$ in (1.2), and this type of convergence is known as Smoluchowski–Kramers (SK) approximation[3, 4, 16, 30, 7]. Then, they found that, the solution $Z^{\epsilon}$ of (1.1), converges to a non-trivial limit. The reason lies in the fact that, in the context of rough path theory, the Itô lift of $X^{\epsilon}$ does not converge to the Itô lift of $X$.
However, in the physical model, it commonly happens that the friction coefficient $M$ is not a constant, and there is a non-linear potential term $F$. In this paper, we explore the limit of (1.1) with the driving signal $X^{\epsilon}$ modeled by a more general stochastic differential equation
$\epsilon^{2}\ddot{X}^{\epsilon}_{t}+M(X^{\epsilon}_{t})\dot{X}^{\epsilon}_{t}=F(X^{\epsilon}_{t})+\dot{W}_{t},\kern 5.0ptX^{\epsilon}_{0}=0,\kern 5.0pt\dot{X}^{\epsilon}_{0}=0.$ (1.3)
We call it a general physical Brownian motion. Setting the velocity $V^{\epsilon}_{t}=\dot{X}^{\epsilon}_{t}$ and momentum $P^{\epsilon}_{t}=\epsilon^{2}V^{\epsilon}_{t}$ respectively, (1.3) is equivalent to
and
We also introduce $Y^{\epsilon}_{t}:=\frac{1}{\epsilon}P^{\epsilon}_{t}$ [7, Page 7944], so that
The system (1.6) has a form of slow-fast system [6], making it possible to find its limit by averaging method. We always assume that equations (1.3) and (1.1) are defined on a complete stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},P)$ with $0<T<\infty$ fixed, and $W$ is a $d$-dimensional standard Brownian motion. Hypothesis on other coefficients are detailed in Section 2. We also remark that, since $X^{\epsilon}$ is of finite variation which can be seen from (1.4) , (1.1) is in fact a random ODE.
The universal limit theorem (also known as continuity of Itô–Lyons map) in rough path theory [19] indicates that, in order to identify the limit of (1.1), it is necessary to find the limit of the lift $\mathbf{X}^{\epsilon}=(X^{\epsilon},\mathbb{X}^{\epsilon})$ of $X^{\epsilon}$, detailed in Section 2. To this end, we will firstly calculate the limit of $X^{\epsilon}$. In contrast to the constant friction coefficient case, we cannot obtain the limit by simply dropping the small mass term $\epsilon^{2}\ddot{X}^{\epsilon}$ in (1.3). In fact, Hottovy et al.[12] showed that, under certain assumptions,
@@
$M(x)J(x)+J(x)M^{\top}(x)=\mathrm{I}_{d}.$ (1.9)
Here $\mathrm{I}_{d}$ is the identity matrix in $\mathbb{R}^{d\times d}.$ For more research in SK approximation in position-dependent friction case, we refer to [5, 11, 17, 28, 27]. As we see in (1.6), SK approximation is naturally linked to stochastic averaging, initiated by Khaminskii [13]. Numerous work concerning about stochastic averaging appeared in recent years, see e.g. [14, 29, 32, 31]. There is also research about averaging principle in rough path framework, see for example in [15, 25]. Therefore, unlike the existing literature, we use a stochastic averaging method to obtain the limit in a much stronger sense. The next step is to identify the limit of $\mathbb{X}^{\epsilon}$, which is more delicate. After these work, we are able to find the effective approximation of (1.1) by utilizing universal limit theorem.
Our current article differs from these existing literature in at least three aspects. The first is that, our main theorems show the convergence of the lifted process $\mathbf{X}^{\epsilon}$ of $X^{\epsilon}$ in the sense of $p$-th moment of $\alpha$-Hölder inhomogeneous rough path distance with $1/3<\alpha<1/2$. This is much stronger than [12] and [27], in which they showed the convergence of the original process $X^{\epsilon}$ in $L^{2}(\Omega;C([0,T];\mathbb{R}^{d}))$ and convergence in distribution, respectively. The second is that, in [12], they essentially assume the tightness of $(X^{\epsilon})_{0<\epsilon\leq 1}$ in $C([0,T];\mathbb{R}^{d}).$ In our paper, we drop this assumption, and obtain a much stronger result, that is, $(\mathbf{X}^{\epsilon})_{0<\epsilon\leq 1}$ is tight in $\mathscr{C}^{0,\alpha}([0,T];\mathbb{R}^{d})$, the closure of the smooth rough paths in $\alpha$-Hölder rough path space $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$, which is a corollary of the boundedness of $(\mathbf{X}^{\epsilon})_{0<\epsilon\leq 1}$ in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$. In [7], they showed the boundedness by utilizing the Gaussian property of $V^{\epsilon}$ in the constant friction case. However, in our setting, $V^{\epsilon}$ is non-Gaussian due to the fact that the friction $M$ depends on the position, which makes the proof more challenging. The third is that, as a consequence of the variable friction again, $V^{\epsilon}$ is non-Markovian, so we cannot pass the limit by a direct application of ergodic theorem when calculating the limit of $\mathbb{X}^{\epsilon}$ as in [7]. To fix this, we prove a strong averaging principle for a specific slow-fast system. We will see in Section 4 that our slow system is only locally Lipschitz, which does not satisfy the commonly assumed condition in the aforementioned reference. Thanks to the structure of our slow system, we employ Poisson equation technique[21, 22, 23] to show the convergence in $L^{p}$ sense.
@@
The rest of this paper is organized as follows. In Section 2, we recall some basic rough path theory, list some notations, and present our main results. In Section 3, we give moment estimate in rough path norm. In Section 4, we display an averaging principle for specific slow-fast system, which is used to pass the limit. After these preparation, we prove the main theorems in Section 5.
2 Preliminary and Main Results
@@
We firstly recall some basic definitions of rough path theory [8, e.g.]. For systematic display of rough path theory, we refer to [18, 20, 9, 8].
Let $X:[0,T]\to\mathbb{R}^{d},$ $\mathbb{X}:\Delta_{[0,T]}\to\mathbb{R}^{d\times d}$, where
$\Delta_{[0,T]}:=\{(s,t)\in[0,T]^{2}:s\leq t\}.$
@@
(i) $X||_{\alpha}:=\sup_{0\leq s\leq t\leq T}\limits\frac{|X_{t}-X_{s}|}{|t-s|^{\alpha}}<\infty,$
(ii) $\mathbb{X}||_{2\alpha}:=\sup_{0\leq s\leq t\leq T}\limits\frac{|\mathbb{X}_{s,t}|}{|t-s|^{2\alpha}}<\infty,$
(iii) (Chen relation) $\mathbb{X}_{s,u}=\mathbb{X}_{s,t}+\mathbb{X}_{t,u}+X_{s,t}\otimes X_{t,u},\kern 5.0pt0\leq s\leq t\leq u\leq T,$
where $X_{s,t}:=X_{t}-X_{s}$ for $(s,t)\in\Delta_{[0,T]}.$
The set of all $\alpha$-Hölder rough path over $\mathbb{R}^{d}$ is denoted by $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}).$ Note that this space is not a linear space due to the nonlinear Chen relation. Given $\mathbf{X},\mathbf{Y}\in\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d}),$ the inhomogeneous $\alpha$-Hölder rough path metric is defined as
@@
This metric is inhomogeneous with respect to the dilation operator $\delta_{\lambda}$ ($\lambda\in\mathbb{R}$), defined by $\delta_{\lambda}((a,b)):=(\lambda a,\lambda^{2}b),\kern 5.0pta\in\mathbb{R}^{d},b\in\mathbb{R}^{d\times d}.$ Besides, the $\alpha$-Hölder homogeneous rough path norm is defined by
[…]
Next we recall concepts of rough differential equations (RDE). We adopt the definition in [8, Section 8.7]. Let $b\in C^{3}_{b}(\mathbb{R}^{e};\mathbb{R}^{e})$, $\sigma\in C^{3}_{b}(\mathbb{R}^{e};\mathbb{R}^{e\times d})$, that is, they are $n$ times continuously differentiable with bounded $k$-th derivative, $k=1,2,3.$ By a solution of the RDE
$dZ_{t}=b(Z_{t})dt+\sigma(Z_{t})d\mathbf{X}_{t},\quad Z_{0}=\xi,$ (2.1)
we mean that
$Z_{t}-Z_{s}=b(Z_{s})(t-s)+\sigma(Z_{s})X_{s,t}+D\sigma(Z_{s})\sigma(Z_{s})\mathbb{X}_{s,t}+R_{s,t},\quad(s,t)\in\Delta_{[0,T]},$
@@
$\lim\limits_{|\mathcal{P}|\to 0}\sup\limits_{\mathcal{P}\subset[0,T]}\sum\limits_{[s,t]\in\mathcal{P}}|R_{s,t}|=0,$
here $\mathcal{P}\subset[0,T]$ means $\mathcal{P}$ is a partition of $[0,T]$, and $|\mathcal{P}|$ is the mesh of $\mathcal{P}$. It turns out that, the Itô–Lyons map of (2.1) is locally Lipschitz with respect to $\rho_{\alpha}$[8, Thereom 8.5]. To be precise, if $\tilde{Z}$ is the solution of
$d\tilde{Z}_{t}=b(\tilde{Z}_{t})dt+\sigma(\tilde{Z}_{t})d\tilde{\mathbf{X}}_{t},\quad\tilde{Z}_{0}=\tilde{\xi}$, (2.2)
[…]
With preliminaries above, we are ready to display our main results. For each $0<\epsilon\leq 1,$ define the second-order process by
$\mathbb{X}^{\epsilon}_{s,t}:=\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u},\quad(s,t)\in\Delta_{[0,T]}.$ (2.3)
Since $X^{\epsilon}$ is of finite variation, the integral above is in the sense of Riemann–Stieltjes. Setting $\mathbf{X}^{\epsilon}=(X^{\epsilon},\mathbb{X}^{\epsilon}),$ it is straightforward to check that $\mathbf{X}^{\epsilon}$ satisfies Chen relation, and that $\mathbf{X}^{\epsilon}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
@@
Now we state our main theorems.
@@
By Kolmogorov criterion for rough paths[8, Theorem 3.1], one verifies that $\mathbf{X}$ lies in $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$ for all $\frac{1}{3}<\alpha<\frac{1}{2}.$
An application of continuity of Itô–Lyons map with careful computation, detailed in Section 5, yield the following theorem.
Theorem 2.4.
Let $b\in C^{3}_{b}(\mathbb{R}^{e};\mathbb{R}^{e})$ and $\sigma\in C^{3}_{b}(\mathbb{R}^{e};\mathbb{R}^{e\times d}).$ Consider random ODE
$dZ^{\epsilon}_{t}=b(Z^{\epsilon}_{t})dt+\sigma(Z^{\epsilon}_{t})dX^{\epsilon}_{t},\quad Z^{\epsilon}_{0}=\xi.$
Under assumptions $(\mathbf{A1})$$(\mathbf{A4})$, $Z^{\epsilon}\to Z$ in $C^{\alpha}([0,T];\mathbb{R}^{e})$ in probability with $Z_{0}=\xi$ and
$dZ^{i}_{t}$
$=$ $[b_{i}(Z_{t})+\sum\limits_{j,k,l,k^{\prime}}\partial_{z_{k}}\sigma_{ij}(Z_{t})\sigma_{kl}(Z_{t})J_{lk^{\prime}}(X_{t})M^{-1}_{jk^{\prime}}(X_{t})$
$+\sum_{j}\sigma_{ij}(Z_{t})S_{j}(X_{t})+\sum_{j,k}\sigma_{ij}(Z_{t})M^{-1}_{jk}(X_{t})F_{k}(X_{t})]dt+\sum_{j,k}\sigma_{ij}(Z_{t})M^{-1}_{jk}(X_{t})dW^{k}_{t},$
where $S_{j}$ is defined in (1.8).
@@
$A^{\top}$: Transpose of a matrix or a vector AA.
$\mathrm{I}_{d}$: Identity matrix of $d\times d$.
@@
$C^{k}(U;V)$: The set consists of $k$-th continuously differentiable functions from $U$ to $V$, where $k\in\mathbb{N},$ and $U$, $V$ are finite-dimensional Banach spaces.
@@ 3 Moment Estimates
In this section, we give several moment estimates for system (1.6).
@@
For each $p\geq 2,$ set $\phi(y)=|y|^{p},$ then
$D\phi(y)=p|y|^{p-2}y,\quad D^{2}\phi(y)=p(p-2)|y|^{p-4}yy^{\top}+p|y|^{p-2}\mathrm{I}_{d}.$
Applying Itô formula to $\phi(Y^{\epsilon}_{t})$,
@@
$+\frac{1}{2\epsilon^{2}}\int_{0}^{t}\mathrm{Tr}[p(p-2)|Y^{\epsilon}_{s}|^{p-4}Y^{\epsilon}_{s}(Y^{\epsilon}_{s})^{\top}+p|Y^{\epsilon}_{s}|^{p-2}\mathrm{I}_{d}]ds,$
where $H^{\epsilon}$ is a martingale. Taking expectation, differentiating and applying $(\mathbf{A1})$ and $(\mathbf{A4})$,
$\frac{d}{dt}\mathbb{E}|Y^{\epsilon}_{t}|^{p}\lesssim-\frac{1}{\epsilon^{2}}\mathbb{E}|Y^{\epsilon}_{t}|^{p}+\frac{1}{\epsilon^{2}}\mathbb{E}|Y^{\epsilon}_{t}|^{p-2}.$ (3.2)
@@
We give estimates by dichotomy. From (1.6),
Rearranging,
Case 1: $\epsilon\geq\sqrt{t-s}$.
By (1.6), Hölder inequality, Fubini theorem and Proposition 3.1,
$\mathbb{E}|X^{\epsilon}_{t}-X^{\epsilon}_{s}|^{p}$ $\leq$ $\frac{1}{\epsilon^{p}}\mathbb{E}\Big(\int_{s}^{t}|Y^{\epsilon}_{u}|du\Big)^{p}$ (3.4)
$\leq$ $\frac{1}{\epsilon^{p}}(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|Y^{\epsilon}_{u}|^{p}du$
$\leq$ $\frac{1}{\epsilon^{p}}(t-s)^{p}\leq(t-s)^{p/2}.$
Case 2: $\epsilon<\sqrt{t-s}$.
From (3.3),